OpenStudy (anonymous):

Calculate the density of a liquid in a beaker. The beaker has a mass of 15.2 grams when empty. The beaker plus an unknown liquid has a mass 35 grams. what is the density of the liquid if its volume is 10.00mL.

OpenStudy (aaronq):

total mass - mass of beaker = 35g-15.2g =19.8 g = liquid's mass d=m/v = 19.8g/10mL = 1.98 g/mL